Q. please enter the missing no. 3,4,8,17,33 ans.50 58 66 99 no define
Answer by stanbon(75887)

please enter the missing no. 3,4,8,17,33 ans.50 58 66 99 no define
----
3
3+1^2 = 4
4+2^2 = 8
8+3^2 = 17
17+4^2 = 33
33+5^2 = 58
=====================
Cheers,
Stan H
================